OpenStudy (anonymous):

Help on getting started please? x^6+7x^3-8=0

OpenStudy (anonymous):

Ah, you again :) Your equation is x^6 + 7x^3 - 8 = 0 right.. Let's do a substitution. Let's say that x^3 = y Therefore, you have y^2 + 7y - 8 =0 Now solve this algebraically, break up the terms.

OpenStudy (anonymous):

x to the 3rd power = -8 x^6+7x^3-8=0 Simple and best practice solution for x^6+7x^3-8=0 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for x^6+7x^3-8=0 equation: Simplifying x6 + 7x3 + -8 = 0 Reorder the terms: -8 + 7x3 + x6 = 0 Solving -8 + 7x3 + x6 = 0 Solving for variable 'x'. Factor a trinomial. (-8 + -1x3)(1 + -1x3) = 0

OpenStudy (anonymous):

Never mind, ^^^ the person above me will tell you how

OpenStudy (anonymous):

Breaking up the term, you would get \[y^{2}+8y-y-8\] \[y(y+8)-1(y+8)\] \[(y-1)(y+8)\] Hence \[y=1\] or \[y=-8\] Now, remember that \[x^{3} = y\] so \[x^{3} = -8\] or \[x^{3} = 1\] Solve both of the equations and you have \[x=-2\] or \[x=1\] I hope my answer helped you. If it solves your qeustion be sure to mark it as Best Response and close the question. Glad to be of help. -swas

OpenStudy (anonymous):

LOL @Jennaluv Equation solver?

OpenStudy (anonymous):

Yes, thought i would try something new....... cx

OpenStudy (anonymous):

But how did you know?

OpenStudy (anonymous):

@Jennaluv : You forgot to edit your answer :) "If it's not what You are looking for type in the equation solver your own equation and let us solve it."